Consider a collection of envelopes consisting of 3 red envelopes​, 2 blue envelopes​, 1 green envelope​, and 1 yellow envelope. If one of the envelopes is selected at​ random, determine the probability that a
green or a yellow envelope is selected

Answer by greenestamps(13219)



Number of possible selections = red plus blue plus green plus yellow = 3+2+1+1 = 7.

Number of selections that are green or yellow = green plus yellow = 1+1 = 2.

P(green or yellow) = 2/7.